OpenStudy (anonymous):

A boat on the ocean is 3 mi from the nearest point on a straight shoreline; that point is 14 mi from a restaurant on the shore. A woman plans to row the boat straight to a point on the short and then walk along the short to the restuarant. If she walks at 3 mi/hr and rows at 2mi/hr, at which point on the shore should she land to minimize the total travel time? Let x be the distance between the nearest point on shore and the point she lands on shore. If T is the time it takes her to get to the restaurant, what is the objective function?

OpenStudy (anonymous):

I am so stuck

OpenStudy (dumbcow):

 objective function: \[T = \frac{\sqrt{x^2+9}}{2} + \frac{14 -x}{3}\]